# Eye-tracked Virtual Reality: A Comprehensive Survey on Methods and Privacy Challenges

### v2

@@ IV-B1 Interaction in VR
Similar to the aforementioned eye-based target selection techniques, another way of interaction is through text input. However, the nature and requirements of the task in text entry are different, and one should consider them in the context of interaction in VR. In addition, gaze information may not only help in interaction, but also indicate potential errors during interaction experience [178].
To this end, He et al. proposed TapGazer [149], which utilizes finger tapping to type on a virtual keyboard by incorporating gaze assistance for word selection. Lu et al. [150] presented BlinkType and NeckType text entry techniques, utilizing blink and neck movements in character selection, respectively. The authors found that using a blink signal is more favorable and reaches the highest word-per-minute rate compared to NeckType and dwell-based text-entry techniques. Another approach introduced by Ma et al. [151] combined brain-computer interfaces and eye-tracking data, which allowed users to compose ten words per minute and achieved a 270 bits per second information transfer rate. Rajanna and Hansen [152] conducted a comparative investigation on applying dwell-time and click inputs in conjunction with gaze information as a confirmation mechanism. The authors also analyzed various keyboard visualization options and found that the most convenient combination was click actions with visualization of the entire keyboard. In addition to these works, Mardanbegi and Thies [179] presented an eye tracking-based interaction toolkit EyeMRTK supporting text entry in Unity.
@@
Gaze-based interaction methods may suffer from inaccuracy and stability issues due to the imprecise detection of target objects by factors such as the accuracy of eye tracking, virtual environment complexity, or the nature of the task. In the following, we discuss eye gaze-based disambiguation techniques addressing these challenges.
@@
Interaction methods that rely solely on the intersection of the gaze ray may cause less accurate detection of targeted objects, especially in scenes with objects at multiple depth levels. To address this issue, depth estimation methods based on the eye have been proposed to improve the interaction quality. Most existing approaches to this problem rely on the relationship between target depth and the vestibulo-ocular reflex (VOR) to fix gaze by moving the eye in the opposite direction of the head. Mardanbegi et al. [153] proposed a VOR-based depth estimation approach using VOR gain obtained by observing pupil centers, as a measure of in-depth computation, rather than the more conventional vergence measure. The VOR gain-based model, which only requires one eye, was evaluated using eye-tracking data from 10 participants. Despite the limited number of participants, the findings revealed that the performance level is comparable with the vergence measure conventionally used in in-depth estimation. Mardanbegi et al. [155] introduced a VOR-based technique to overcome disambiguation problems during gaze interaction by jointly utilizing eye and head movements to estimate gaze depth, leading to a better performance compared to conventional vergence-based methods in gaze depth estimation for the virtual objects located deep in the scene.
@@
In another work, Weier et al. [154] built a regression model to improve gaze depth estimation by combining vergence measures with other features acquired through conventional ray-tracing techniques based on the users’ point of regard. The model yielded remarkable accuracy regarding the average deviation from the benchmark depth over a 6-meter range. At the same time, vergence measures offered accurate prediction only up to a meter. Similarly, disambiguation is also a significant challenge in hands-free interaction techniques. Chen et al. [156] evaluated disambiguation techniques such as head gaze, speech, and foot tap in different timing scenarios and found that head gaze outperformed other methods in eliminating disambiguation.
@@
User intent information is also helpful in addressing disambiguation issues that arise in VR interaction based on eye tracking. By predicting user intent, it is possible to identify forthcoming movements of users and objects with which users will interact. To this end, several works utilized eye-tracking-based approaches. For example, Keshava et al. [157] demonstrated the capability of point-of-regard regions to predict four different alignment tasks with cubes, achieving an $F_{1}$-score of $0.51$. David-John et al. [158] employed eye-tracking features to train a logistic regression model and revealed that characteristics of focal attention, gaze velocity, fixation, and saccade dynamics are highly relevant for predicting user intention. Furthermore, Hu et al. [180] proposed EHTask, a novel task classification method utilizing eye and head movement features, and showed that eye-tracking features differ based on the tasks in VR. In addition, EHTask performed better than state-of-the-art methods in task classification in 2D viewing. Similarly, for intent prediction, Alghofaili et al. [159] introduced an approach for predicting navigation assistance needs based on eye gaze information and long short-term memory networks (LSTMs), approaching a prediction accuracy of over $80\%$.
@@
Hand redirection methods are essential for addressing the discrepancies between hand movements in virtual and real environments, as perfect alignment between the two is not always possible. Discrepancies can lead to difficulties during hand interactions, such as the inability to interact with the intended objects. Eye-tracking-based methods can assist in identifying the user’s intent and direct the hand to the correct location.
Cheng et al. [160] utilized a gaze-based hand redirection method to bridge the gap between real and virtual worlds with their proposed Sparse Haptic Proxy framework. This framework consists of geometric primitives that mimic the haptic feedback of virtual objects. The authors employed the haptic re-targeting technique [181] that utilizes users’ eye and hand behaviors to predict their intention and redirect their hand to a physical proxy. User studies showed that the proposed method achieved an accuracy of 97.5% for gaze-based user intention prediction and reported the maximal acceptable hand redirection angle as $40^{\circ}$. Zenner et al. [162] proposed Blink-Suppressed Hand Redirection as a hand redirection approach based on a body wrapping algorithm [160]. Unnoticeable instant changes are performed by exploiting the natural blink motion of the eye, and real and virtual hand offset is adjusted with slight modifications when the eyes are open.
REACH+ of Gonzalez et al. [161] combines eye-tracking data and hand motion to determine user intention in redirection frameworks. The method aims to overcome challenges through physical interaction in virtual spaces. REACH+ redirects the users’ hands to the intended target within arm-reachable range, improving the feeling of realism. In recent work, Sendhilnathan et al. [164] accomplished error detection in gesture input with the help of gaze dynamics. The authors analyzed the gaze patterns following gesture input and classified the input into three types: correct input, input recognition errors, and user errors. Utilizing only gaze features such as fixation duration, saccade amplitude, and gaze velocity, the researchers successfully achieved a classification area under the curve of a receiver operating characteristic curve for one-vs-rest score (AUC-ROC-OVR) of 0.78 with a temporal convolutional network (TCN) and examined the consistency of such classification across tasks. More in terms of object manipulation tasks, Yu et al. [163] designed an interaction technique combining hand gestures with gaze information for object manipulation tasks in VR. The authors evaluated four alternative combination techniques while varying the object distance from the agent in the virtual environment. According to user studies, the authors showed that combining hand and gaze improved the usability and efficiency of object manipulation tasks in large environments where objects are not within the arm’s reach distance.
@@
Another challenge in VR interaction, especially during mobile activities, is the limited space in the real world. Nguyen and Kunz [165] proposed a discrete rotation technique to minimize the required area in a real workspace. The authors first assessed the threshold for the discrete rotation in virtual scenes and obtained $9.1^{\circ}$ and $2.4^{\circ}$ for closed and open eyes, respectively. Then, they applied discrete scene rotation while the agent was walking and reduced the required area by 20%. An alternative method to facilitate virtual walking activity in constrained physical space was proposed by Sun et al. [182]. The authors employed a path-planning algorithm that considered positions of pre-existing static and dynamic obstacles, including walls and furniture, and multiple VR users within a designated room. Furthermore, the algorithm used the saccadic suppression periods to facilitate efficient path planning, and the authors presented a subtle gaze direction technique to increase the number of saccades, considerably enhancing the redirection gain. From a user perception perspective, Brument et al. [183] studied the influence of translational and rotational motion in the context of rotation gains and found, with the gaze analyses, that rotation gains can alter the user behavior, which may be important when implementing the redirection techniques.
Predicting future physical positions can help optimize the use of limited physical area [166, 167]. Stein et al. [166] proposed to use future location predictions obtained through an LSTM model utilizing eye tracking, and demonstrated that their model can predict users’ position 2.5 seconds in advance, with an average error of 65 cm. A similar approach relying on the LSTM model was proposed by Bremer et al. [167], evaluating short-term (i.e., 50 ms) and long-term (i.e., 2.5 seconds) predictions separately. They achieved the optimal performance in short-term predictions by utilizing position and orientation features. In contrast, adding eye-tracking features resulted in the best performance for long-term predictions. The authors also stated that short-term predictions might be valuable for optimizing computational resources for rendering or transmitting bandwidth in streaming activities.
@@ IV-B2 Predictive Gaze Analysis
Predicting the future gaze locations enhances the efficiency of gaze-based rendering techniques, particularly foveated rendering. Therefore, methods combining gaze prediction and foveated rendering have gained strong momentum in recent years. This section discusses the algorithms for gaze location prediction, scanpath prediction, and saliency map generation. We provide an overview of the papers from this section in Table III.
TABLE III: A non-exhaustive overview of papers discussed in Section IV-B2.
@@
Gaze prediction [169] DGaze: CNN model using object-, head-, and saliency-based features
@@
Foveated rendering methods depend on accurate gaze prediction, which directly impacts the performance and usability of foveated rendering applications. Therefore, researchers have dedicated significant attention to exploring and implementing effective gaze prediction techniques, leveraging various methods from statistical approaches to neural networks. For instance, Deep Future Gaze (DFG), proposed by Zhang et al. [187], is a GAN-based model that exploits a spatial-temporal CNN as an encoder and anticipates upcoming gaze positions while generating future frames based on the current ones. A by-product of their work is an object search task (OST) dataset, considered one of the most extensive egocentric datasets. Hu et al. [184] proposed SGaze that utilizes eye-gaze information collected by an integrated eye tracker to forecast future gaze locations. SGaze was designed as a data-driven model based on eye-head coordination for real-time gaze prediction and is based on statistical models exploiting the relation of gaze and head movements, without any additional hardware or eye-tracker support. This model took the latency in the eye and head movements into account and achieved a better angular distance value than the baseline, which used the screen center as the gaze point.
More recently, Hu et al. [169] also presented DGaze as a CNN-based gaze prediction algorithm incorporating dynamic object positions, head movements, and saliency features extracted with SAM-ResNet saliency predictor [199]. Object positions, head velocity data, and previous gaze location sequences are processed with a sequence encoder model to increase the precision. This technique was applied in real-time, and near-future gaze predictions corresponded to short-term predictions such as 200-1000 ms. DGaze achieved better angular distance values in dynamic and stationary scenes than the prior method, SGaze [184].
@@
FixationNet, proposed by Hu et al. [185], is a neural network utilized to predict near-future gaze locations up to 600 ms in task-oriented virtual experiences utilizing eye-tracking data, head movements, saliency maps, and task-related information. Data from 27 individuals during the visual searching task in VR indicated a strong correlation between fixation positions and head movements, saliency maps, and content-related information. Besides estimating the gaze position, the proposed model outperforms state-of-the-art models in gaze prediction tasks for 150-600 ms in task-oriented and free-viewing experiments. Another gaze prediction method, by Koulieris et al. [186], focuses on heavily task-oriented games. The proposed approach exploits the correlation between the game variables in their current states and gaze locations. Moreover, predicted gaze locations were used in the dynamic disparity manipulation method to improve the depth sense in the game scenes. Vielhaben et al. [188] predicted the users’ future head orientation using past head movements and eye-tracking data by training a linear support vector machine (SVM). Their model can predict future viewports without utilizing any information related to the content in spherical videos; hence, it can help improve rendering performance and reduce the transmission load.
@@
To provide a comprehensive overview of gaze prediction, we discuss the scanpath prediction, referring to the sequence of gaze points in a given virtual environment. Unlike gaze prediction techniques that forecast individual samples, scanpath prediction involves predicting the sequence of gaze points, allowing for the estimation of complete human attention on a given scene. While these two concepts are related, they serve different purposes and require different types of models and algorithms for prediction.
To this end, Assens et al. [189] proposed SaltiNet, a deep CNN trained first to generate saliency volumes that capture temporal information, and then the scanpath is sampled from the saliency volumes. In another work, Assens et al. [190] designed a novel GAN model, called PathGAN, that employs a convolutional-recurrent architecture for both the encoder and decoder to predict scanpath. Recently, Martin et al. [191] introduced another GAN-based scanpath prediction model, ScanGAN360, which deploys a novel GAN objective function based on dynamic time warping. Unlike the ScanGAN360, Zhu et al. [192] suggested generating scanpaths by building a graph from a saliency map. The saliency map is first binarized and clustered into centers. Then, a weighted graph is created with the cluster centers as nodes, and a scanpath can be generated from the graph.
@@
A saliency map, also known as a heat map or attention map, is a visual representation of eye-gaze data on an image or a scene. It serves as a visualization tool that highlights the regions that attract the most attention from human observers. Figure 7 illustrates a saliency map, which shows the distribution of human attention over an image. The saliency map lays the foundation for numerous gaze-based VR applications and is important in understanding human visual attention.
Different approaches exist in the literature for saliency map estimation and generation. John et al. [193] benchmarked four different methods for saliency map generation in VR. The authors suggested utilizing the modified Gaussian kernel [200] with a scale of 5% for $360^{\circ}$ saliency map generation due to ease of implementation. Pfeiffer and Memili [194] presented an approach that tackles challenges such as changing perspectives, dynamically moving objects, and depth of fixations in the generation of saliency maps for 3D environments. The authors aggregated the gaze data from multiple users and used textures to represent visual attention distribution at the object level. Their method offers high-quality saliency maps for multi-perspective eye-tracking recordings.
@@
Sitzmann et al. [195] studied the difference in saliency patterns between desktop and stereoscopic vision among 169 participants and adapted traditional saliency prediction methods according to user-specific behaviors such as particular fixation biases. Nguyen et al. [196] researched predicting head movement using an LSTM model in conjunction with saliency maps. Maranes et al. [197] exploited a dataset of 3259 users watching cinematographic VR movies and measured user attention via saliency maps. Closely related to visual saliency, Ma et al. [198] aimed at salient object detection (SOD) in $360^{\circ}$ panorama images with the help of eye tracking and proposed a novel dataset, i.e., 360-SSOD, including manually annotated object-level saliency ground truth with balanced semantic distribution compared to existing datasets.

### v3

@@ IV-B1 Interaction in VR
Similar to eye-based target-selection techniques, another form of interaction is text input. However, the nature and requirements of text-entry tasks differ and should be considered in the context of VR interactions. Moreover, gaze information can support not only interaction itself but also reveal potential errors [153].
To this end, [126] proposed TapGazer, utilizing finger tapping to type on a virtual keyboard by incorporating gaze assistance for word selection. [127] presented BlinkType and NeckType text entry techniques, utilizing blinks and neck movements in character selection, respectively, and found that using blinks is more favorable and reaches the highest word-per-minute rate compared to NeckType and dwell-based text-entry techniques. Another approach [128] combined brain-computer interfaces and eye tracking, enabling users to compose ten words per minute and achieving an information transfer rate of 270 bits per second. [129] conducted a comparative investigation on applying dwell-time and click inputs in conjunction with gaze as a confirmation mechanism, also analyzing various keyboard visualization options, and found that the most convenient combination was click actions with the entire keyboard. Additionally, [154] presented EyeMRTK, an eye-tracking-based toolkit that supports text entry in Unity.
@@
Gaze-based interaction methods may suffer from inaccuracy and stability issues due to the imprecise detection of target objects, which can be influenced by factors such as inaccurate eye tracking or the complexity of the virtual environment. We discuss eye-gaze-based disambiguation techniques that address these challenges.
@@
Interaction methods that rely solely on the intersection of the gaze ray may cause less accurate detection of targeted objects, especially in scenes with objects at multiple depth levels. To address this issue, depth estimation methods based on the eye have been proposed to improve the interaction quality. Most approaches rely on the relationship between target depth and the vestibulo-ocular reflex (VOR) to fix gaze by moving the eye in the opposite direction of the head. [130] proposed a VOR-based depth estimation approach using VOR gain obtained by observing pupil centers, as a measure of in-depth computation, rather than the more conventional vergence measure. The VOR gain-based model, which only requires one eye, revealed that its performance is comparable to that of the vergence measure conventionally used in in-depth estimation. [132] introduced a VOR-based technique to overcome disambiguation problems during gaze interaction by jointly utilizing eye and head movements to estimate gaze depth, leading to a better performance compared to conventional vergence-based methods in gaze depth estimation for the virtual objects located deep in the scene.
@@
In another work, [131] developed a regression model to enhance gaze depth estimation by integrating vergence measures with other features obtained through conventional ray-tracing techniques based on the users’ point of regard. The model yielded remarkable accuracy regarding the average deviation from the benchmark depth over a six-meter range, while vergence measures offered accurate prediction only up to a meter. Similarly, disambiguation is also a significant challenge in hands-free interaction. [133] evaluated disambiguation techniques such as head gaze, speech, and foot tap in different timing scenarios and found that head gaze outperformed other methods in eliminating disambiguation.
@@
User intent information is also helpful in addressing disambiguation issues that arise in VR interaction based on eye tracking. By predicting user intent, it is possible to identify forthcoming movements of users and objects with which users will interact, and several works have utilized eye tracking for this. For example, [134] demonstrated the capability of point-of-regard regions to predict four different alignment tasks with cubes, achieving an $F_{1}$-score of $0.51$. [135] employed eye-tracking features to train a logistic regression model and revealed that characteristics of focal attention, gaze velocity, and scanpath dynamics are highly relevant for predicting user intention. Furthermore, [155] proposed EHTask, a novel task classification method that utilizes eye and head movement features, and demonstrated that eye-tracking features differ based on the tasks in VR. Additionally, EHTask outperformed state-of-the-art methods in task classification for 2D viewing. Similarly, for intent prediction, [136] introduced an approach that predicts navigation assistance needs based on eye gaze information and long short-term memory networks (LSTMs), achieving a prediction accuracy of over 80%.
@@
Hand redirection methods are essential for addressing the discrepancies between hand movements in virtual and real environments, as perfect alignment between the two is not always possible. Discrepancies can lead to difficulties during hand interactions, such as the inability to interact with the intended objects. Eye-tracking-based methods can help identify the user’s intent and direct the hand to the correct location.
To this end, [137] utilized a gaze-based hand redirection method to bridge the gap between real and virtual worlds with their framework, Sparse Haptic Proxy. This framework consists of geometric primitives that mimic the haptic feedback of virtual objects. The authors employed the haptic re-targeting technique [156] that utilizes users’ eye and hand behaviors to predict intentions and redirect their hand to a physical proxy, and showed that it achieved an accuracy of 97.5% for gaze-based user intention prediction, while reporting the maximal acceptable hand redirection angle as $40^{\circ}$. [139] proposed Blink-Suppressed Hand Redirection as a hand redirection approach based on a body wrapping algorithm [137]. Unnoticeable instant changes are performed by exploiting the natural blink motions, and the real and virtual hand offset is adjusted with slight modifications when the eyes are open.
REACH+ [138] combines eye-tracking data and hand motion to determine user intention in redirection frameworks, and aims to overcome challenges through physical interaction in virtual spaces. REACH+ redirects the users’ hands to the intended target within arm-reachable range, enhancing the sense of realism. In recent work [141], error detection in gesture input was accomplished with the help of gaze dynamics, and the gaze patterns following gesture input were analyzed to classify the input into three types: correct input, input recognition errors, and user errors. Utilizing only gaze features such as fixation duration, saccade amplitude, and gaze velocity led to a classification area under the curve of a receiver operating characteristic curve for one-vs-rest score (AUC-ROC-OVR) of 0.78 with a temporal convolutional network (TCN). Considering object manipulation tasks, [140] designed an interaction technique combining hand gestures with gaze information for object manipulation tasks in VR, and evaluated four alternative combination techniques while varying the object distance from the agent in the virtual environment. Combining hand and gaze improved the usability and efficiency of object manipulation tasks in large environments where objects are not within arm’s reach.
@@
Another challenge in VR interaction, especially during mobile activities, is the limited space in the real world. [142] proposed a discrete rotation technique to minimize the required area in a real workspace. They first assessed the threshold for the discrete rotation in virtual scenes and obtained $9.1^{\circ}$ and $2.4^{\circ}$ for closed and open eyes, respectively. Then, they applied discrete scene rotation while the agent was walking, reducing the required area by 20%. An alternative method to facilitate virtual walking activity in constrained physical spaces employed a path-planning algorithm that considered the positions of pre-existing static and dynamic obstacles, including walls and furniture, as well as multiple VR users within a designated room [157]. Furthermore, the algorithm utilized saccadic suppression periods to facilitate efficient path planning, and the authors presented a subtle gaze direction technique to increase the number of saccades, thereby considerably enhancing the redirection gain. From a user perception perspective, [158] studied the influence of translational and rotational motion in the context of rotation gains and found, through gaze analyses, that rotation gains can alter user behavior, which may be important when implementing redirection techniques.
Predicting future physical positions can help optimize the use of limited physical area [143, 144]. [143] proposed to use future location predictions obtained through an LSTM model utilizing eye tracking, and demonstrated that their model can predict users’ position 2.5 seconds in advance, with an average error of 65 cm. A similar approach, relying on the LSTM model and evaluating short-term (i.e., 50 ms) and long-term (i.e., 2.5 seconds) predictions separately, achieved optimal performance in short-term predictions by utilizing position and orientation features [144]. In contrast, adding eye-tracking features resulted in the best performance for long-term predictions. The authors stated that short-term predictions might be valuable for optimizing computational resources for rendering or transmitting bandwidth in streaming activities.
@@ IV-B2 Predictive Gaze Analysis
Predicting the future gaze locations enhances the efficiency of gaze-based rendering techniques, particularly foveated rendering. Therefore, methods that combine gaze prediction and foveated rendering have gained significant momentum in recent years. This section discusses the algorithms for gaze location prediction, scanpath prediction, and saliency map generation. Table III provides an overview of the papers in this section.
TABLE III: A broad overview of papers in Section IV-B2.
@@
Gaze prediction [146] DGaze: CNNs with object, head, and saliency data
@@
Foveated rendering methods depend on accurate gaze prediction, which directly impacts the performance and usability of foveated rendering applications. Therefore, researchers have devoted significant attention to exploring and implementing effective gaze prediction techniques, leveraging various methods, ranging from statistical approaches to neural networks. For instance, Deep Future Gaze (DFG) [162] is a GAN-based model that exploits a spatial-temporal CNN as an encoder and anticipates upcoming gaze positions while generating future frames based on the current ones. A by-product of their work is an object search task (OST) dataset, considered one of the most extensive egocentric datasets. [159] proposed SGaze that utilizes eye-gaze information collected by an integrated eye tracker to forecast future gaze locations. SGaze was designed as a data-driven model based on eye-head coordination for real-time gaze prediction, utilizing statistical models that exploit the relationship between gaze and head movements, without requiring any additional hardware or eye-tracker support. This model took into account the latency in eye and head movements, achieving a better angular distance value than the baseline, which used the screen center as the gaze point. More recently, [146] presented DGaze as a CNN-based gaze prediction algorithm incorporating dynamic object positions, head movements, and saliency features. Object positions, head velocity data, and previous gaze location sequences are processed using a sequence encoder model to enhance precision. This technique was applied in real-time, and near-future gaze predictions corresponded to short-term predictions, such as those within 200-1000 ms. DGaze achieved better angular distances in dynamic and stationary scenes than SGaze [159].
@@
FixationNet is a neural network utilized to predict near-future gaze locations up to 600 ms in task-oriented virtual experiences, using eye-tracking data, head movements, saliency maps, and task-related information [160]. Evaluations during the visual searching task in VR indicated a strong correlation between fixation positions and head movements, saliency maps, and content-related information. Besides estimating the gaze position, the proposed model outperforms state-of-the-art models in gaze prediction tasks for 150-600 ms in task-oriented and free-viewing experiments. Another gaze prediction method focuses heavily on task-oriented games and exploits the correlation between the game variables in their current states and gaze locations [161]. Moreover, predicted gaze locations were utilized in the dynamic disparity manipulation method to enhance depth perception in game scenes. [163] predicted the users’ future head orientation using past head movements and eye-tracking data with a support vector machine (SVM). This model can predict future viewports without utilizing any information related to the content in spherical videos; hence, it can help improve rendering performance and reduce the transmission load.
@@
Unlike gaze prediction techniques that forecast individual samples, scanpath prediction involves predicting the sequence of gaze points, enabling the estimation of complete human attention to a given scene. While these concepts are related, they serve different purposes and require different models and algorithms for prediction.
To this end, [164] proposed SaltiNet, a deep CNN trained first to generate saliency volumes that capture temporal information, and then the scanpath is sampled from the saliency volumes. In another work, [165] designed a novel GAN model, PathGAN, that employs a convolutional-recurrent architecture for both the encoder and decoder to predict scanpath. Recently, [166] introduced another GAN-based scanpath prediction model, ScanGAN360, which deploys a novel GAN objective function based on dynamic time warping. Unlike the ScanGAN360 [167], which suggested generating scanpaths by building a graph from a saliency map, the saliency map is first binarized and clustered into centers. Then, a weighted graph is created with the cluster centers as nodes, and a scanpath can be generated from the graph.
@@
A saliency map, also known as a heat map or attention map, is a visual representation of eye-gaze data on an image or a scene. It serves as a visualization that highlights the regions that attract the most attention from humans. Figure 6 illustrates a saliency map, showing the distribution of human attention. Saliency maps lay the foundation for numerous gaze-based VR applications.
@@
Several approaches exist in the literature for estimating and generating saliency maps. [168] benchmarked four different methods for saliency map generation in VR, and suggested utilizing the modified Gaussian kernel [174] with a scale of 5% for $360^{\circ}$ saliency map generation due to ease of implementation. [169]’s approach addresses challenges such as changing perspectives, dynamically moving objects, and depth of fixations in generating saliency maps for 3D environments. The authors aggregated user gaze data and utilized textures to represent the distribution of visual attention at the object level. Their method offers high-quality saliency maps for multi-perspective eye-tracking recordings.
[170] investigated the difference in saliency patterns between desktop and stereoscopic vision and adapted traditional saliency prediction methods according to user-specific behaviors such as particular fixation biases. [171] researched predicting head movement using an LSTM model in conjunction with saliency maps. [172] utilized a dataset of 3,259 users watching cinematographic VR movies and measured user attention using saliency maps. Closely related to visual saliency, [173] aimed at salient object detection (SOD) in $360^{\circ}$ panorama images with eye tracking and proposed a novel dataset, 360-SSOD, including object-level saliency ground truth with balanced semantic distribution compared to existing datasets.